OpenStudy (anonymous):

Which set of side lengths will form a right triangle? 5, 11, 13 9, 24, 25 7, 24, 25 8, 13, 17

Parth (parthkohli):

The sides are given in the form \((a,b,c)\). The only step is to be performed is to check if \(a^2 + b^2 = c^2\)

OpenStudy (unklerhaukus):

the sums of the squares of the smaller side lengths must equal the square of the large side for it to be a right angles triangle

OpenStudy (anonymous):

Check for third one..

Parth (parthkohli):

As said by the Great Mr. Pythagorean, a^2 + b^2 = c^2 in a right triangle ^_^

OpenStudy (anonymous):

\[\huge \color{blue}{Hypotenuse^2 = Base^2 + Height^2}\]